What kind of experiment (what kind of cross) could you perform to determine the genotype of the selected petunia plant? Explain briefly. b) How will the cross in (a) help you determine the genotype of your red-flowered petunia? That is, explain how will the results from this cross differ if the red-flowered petunia is A/A versus A/a? c) You end up choosing two red plants that are both of the genotype A/a and you cross them. If the cross produces six plants, what is the probability that four of them will be white and two will be red? Show/explain your work for full marks.
Answer #1

A) We should do test cross to determine the genotype of the selected petunia plant. In this cross, we have to cross our unknown genotype plant with homozygous recessive parent.

B)

In this cross, if we found 1:1 ratio of red and white flower plants in the progeny, then the genotype of plant is Aa

When all progenies are of red flower, then the genotype is AA.

C)

Parents: Aa × Aa

Punnet square:

Gametes A a
A AA( red) Aa( red)
a Aa( red) aa( white)

We have to do binomial expansion this problem because order is not specified.

The probability of white plant (P) is 1/4

The probability of red plant is (q) is 3/4

The total number of progeny are (n) = 6

White progeny number (x) = 4

Probability ={ n! /x( n-x)! }× (P) ^x × (q) ^(n-x)

={ 6! /4! (2)! }×(1/4) ^4×(3/4) ^2

=( 30/2) ×(1/256) ×(9/16)

= 15× (1/256) ×(9/16)

= 135/4096

= 0.03295

That's why probability that four of them will be white and two will be red is 0.03295